A car travels 100m as it accelerates from rest at 25m/s. Calculate the acceleration.
1
Expert's answer
2018-09-13T10:35:08-0400
We can find the acceleration from the kinematic equation:
v^2=v_0^2+2as,
here, v_0=0 is the initial velocity of the car, v=25 m⁄s is the final velocity of the car, a is the acceleration of the car and s=100 m is the distance traveled by the car.
Then, from this equation we can find the acceleration of the car:
a=v^2/2s=(25 m/s)^2/(2∙100 m)=3.125 m/s^2 .